OpenStudy (anonymous):

Find the slope of the line. -3y+4x=-27

OpenStudy (ybarrap):

Do you know how to write in y-intercept format?

OpenStudy (ybarrap):

Like this - y=mx+b?

OpenStudy (anonymous):

I'm sorry I had to do something. Yes I know how write the y- intercept form.

OpenStudy (anonymous):

the slope is 4/3

OpenStudy (anonymous):

slope = - x-intercept / y-intercept m = -4 / -3 m=4/3

OpenStudy (anonymous):

The equation for a line is y=mx+b where m is the slope and b is the y intercept, so you just have to reformat the equation to get the answers. I will show you how

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so first you want to move the 4x to the right side of the equation, subtract it from both sides. now you have 3y=-4x-27, then you want to divide both sides (the entire thing) by -3. So that gives you Y= -4/-3x + 27/-3 (a negitive divided by a negative gives you a positive slope) and 27/-3 is -9, so your equation is y=3/4x-9 (your slope is 4/3, and your y intercept is -9) :D hope that helps!

OpenStudy (anonymous):

Yes. Thank You.